OpenStudy (anonymous):

A ball is thrown straight up with an initial velocity of 35 mph. How high does it go (as measured from its initial height)? Answer in meters.

OpenStudy (aaronq):

 1 mile = 1.60934 km = 1609.34 m 35 mph = 56326.9 m/h m/h to m/s 56326.9 m/h \(*\dfrac{1~h}{60~min}*\dfrac{1~min}{60~sec}=15.646361~m/s \approx ~15.7~ m/s\) use: \(\sf v_f^2=v_i^2+2a\Delta y\) the ball comes to a stop at the top, so \(\sf v_f=0\) \(\sf (0)^2=(15.7~ m/s)^2+2(-9.8~/s^2)\Delta y\) \(\Delta y=12.49~m \approx 12.5~m\)